OpenStudy (pflori1234_pop):

Two point charges each with 9.0×10^−2 C of charge are 8.0×10^−2 m apart. Knowing that k is 9×10^9N⋅m^2/C^2 (the proportionality constant for Coulomb's law), find the force between the charges.

OpenStudy (pflori1234_pop):

Coulomb's law: F=K * q1q2/ d2

OpenStudy (pflori1234_pop):

@TheSmartOne

OpenStudy (pflori1234_pop):

Can you help?

TheSmartOne (thesmartone):

q1 and q2 are the charges d is the distance k is a constant they've given you all the information you needed, all you need to do is plug in! :D

OpenStudy (pflori1234_pop):

I'm confused how to even start multiplying tho

TheSmartOne (thesmartone):

Ah, I see. It looks confusing with all the scientific notation and the units. Let's only look at the numerator for now :D (9*10^9N⋅m^2/C^2) * (9.0*10^−2 C) * (9.0*10^−2 C) The first one is K, the second and third one are the two charges

OpenStudy (pflori1234_pop):

so (9*10^9N⋅m^2/C^2) * (81*10^-4c)?

OpenStudy (pflori1234_pop):

I'm still confused idk if thats right

TheSmartOne (thesmartone):

Let me write it out using LaTeX \(\Large\sf \frac{9\cdot10^9N⋅m^2}{C^2} \times \frac{9.0C}{10^2} \times \frac{9.0C}{10^2}\) I used this rule of exponents to get the 10^2 in the denominator \(\Large a^{-b} = \frac{1}{a^b}\)

TheSmartOne (thesmartone):

you got it almost correct except right at the end you should have C^2 so we have now: \(\Large\sf \frac{9\cdot10^9N⋅m^2}{C^2} \times \frac{81.0C^2}{10^4}\)

OpenStudy (pflori1234_pop):

would it be 729*10^9n*c^2*m^2/ 10000c^2

OpenStudy (pflori1234_pop):

i dont understand for my hw it says the answer should be an integer

TheSmartOne (thesmartone):

\(\Large\sf \frac{9\cdot10^9N⋅m^2}{C^2} \times \frac{81C^2}{10^4}\) all you have to do is multiply the numerators with the numerator and the denominator with the denominator basically 9 * 10^9 * 81 = (9 * 81) * 10^9 so don't get scared of that 10^9 (:

TheSmartOne (thesmartone):

let's leave the denominator as 10^4 and the C^2 cancel since we have it in both numerator and denominator and you can divide 10^9/10^4 so we can can get rid of 10^4 from the denominator Hint: \(\Large \frac{a^b}{a^c} = a^{b-c}\)

TheSmartOne (thesmartone):

and currently, we're only simplifying the numerator! We haven't considered the denominator yet!

OpenStudy (pflori1234_pop):

would it be 729*10^9n*c^2*m^2/ c^2 *10^4

TheSmartOne (thesmartone):

yes, but we can simplify that to become  that's just the numerator, however!